OpenStudy (anonymous):

If y varies directly as​ x, and y=12 when x=4​, find y when x=9.

OpenStudy (anonymous):

Direct variation, do you know the formula?

OpenStudy (anonymous):

y=kx

OpenStudy (anonymous):

I am getting the wrong answer and only have one try left for this online course. could you help?

OpenStudy (anonymous):

Substitute the given values.

OpenStudy (jhannybean):

Start by solving for k k = y/x = 12/4 = ?

OpenStudy (anonymous):

i got 3=k and 3=x. says x is wrong

OpenStudy (jhannybean):

One step at a time. k = 3. substitute this back in to your formula. y =3x Now when x = 9 what would y =? y = 3 * 9 = ?

OpenStudy (anonymous):

oh okay. i divided

OpenStudy (jhannybean):

mmhmm.

OpenStudy (anonymous):

thx